OpenStudy (anonymous):

The mean (average) of the two numbers, a and b,is given by the equation. m=a+b/2 Solve this equation for a Answer options: a=2m-2b a=b-2m a=2m/b a=2m-b

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

a=2m-b is what i get off the top of my head

OpenStudy (gw2011):

If the equation is: m=(a+b)/2, then the solution is: 2m=a+b we can rewrite this equation as: a+b=2m a=2m-b

myininaya (myininaya):

m=(a+b)/2 2m=a+b 2m-b=a a=2m-b